OpenStudy (anonymous):

average weight of 3 containers a b and c is 3.2 kg. container a is 2x as heavy as container b. Container b is 400g heavier than container c. find weight of container c. thanks,

OpenStudy (anonymous):

a + b +c -------- = 3.2 kg ---> 1st equation 3 a = 2b ---> 2nd equation b = c + 0.4kg ---> 3rd equation 0.4 kg = 400g b - 0.4kg = c a + b + c --------- = 3.2kg a = 2b, c = b - 0.4kg 3 2b + b +b - 0.4kg ---------------- = 3.2kg 3 4b - 0.4kg = 9.6kg 4b = 10kg 10kg b = ----- 4 b = 2.5kg a = 5kg c = 2.1kg

OpenStudy (anonymous):

kelu_baog had answered it correctly but isn't that 'weight' is also important because \[mass \neq weight\]